OpenStudy (anonymous):

find 3 consecutive integers whose sum is -93

OpenStudy (saifoo.khan):

(x)+(x+1)+(x+2) = -93

OpenStudy (anonymous):

but what are the 3 numbers?

OpenStudy (saifoo.khan):

3x+3 = -93 3x = -96 x = -96 / 3 x = -31

OpenStudy (saifoo.khan):

therefore numbers r -13 -14 -15

OpenStudy (anonymous):

okay, thank you.

OpenStudy (saifoo.khan):

:)

OpenStudy (anonymous):

lol genius

OpenStudy (saifoo.khan):

narayain!

OpenStudy (anonymous):

saifoo a ques for u

OpenStudy (saifoo.khan):

Sure.

OpenStudy (saifoo.khan):

kya dhond rahe ho?

OpenStudy (anonymous):

Surely numbers are -31, -32 and -33? The sum of 13, 14, and -15 is -42.